An investor purchases one municipal bond and one corporate bond that pay rates of return of 4% and 5.5%, respectively. If the investor is in the 20% tax bracket, his after-tax rates of return on the municipal and corporate bonds would be, respectively, _____.
Answer #1

After-tax rates of return on the municipal bond = 0.04(1 - 0) = 0.04 or 4%

After-tax rates of return on the corporate bond = 0.055(1 - 0.20) = 0.044 or 4.4%